# GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

### v3

@@ Abstract.
Text classification is a fundamental task in data mining, pivotal to various applications such as tabular understanding and recommendation. Although neural network-based models, such as CNN and BERT, have demonstrated remarkable performance in text classification, their effectiveness heavily relies on abundant labeled training data. This dependency makes these models less effective in dynamic few-shot text classification, where labeled data is scarce, and new target labels frequently appear based on application needs. Recently, large language models (LLMs) have shown promise due to their extensive pretraining and contextual understanding ability. Current approaches provide LLMs with text inputs, candidate labels, and additional side information (e.g., descriptions) to classify texts. However, their effectiveness is hindered by the increased input size and the noise introduced through side information processing. To address these limitations, we propose a graph-based online retrieval-augmented generation framework, namely GORAG, for dynamic few-shot text classification. Rather than treating each input independently, GORAG constructs and maintains a weighted graph by extracting side information across all target texts. In this graph, text keywords and labels are represented as nodes, with edges indicating the correlations between them. To model these correlations, GORAG employs an edge weighting mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a minimum-cost spanning tree tailored for each text input. Empirical evaluations demonstrate that GORAG outperforms existing approaches by providing more comprehensive and precise contextual information.
Large language models, online retrieval-augmented generation, few-shot learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/02
Figure 1. An example of the DFSTC task with two rounds. In round 1, an example text is classified as Machine Learning. In round 2, an example text is classified as Computer Vision.
@@ 1. Introduction
Text classification is a fundamental task of text data mining and is connected to various tasks, such as tabular understanding (Chen et al., 2024c; Wang et al., 2024c; Wen et al., 2024a, b; He et al., 2025; Zheng et al., 2023), and recommendation (Du et al., 2024a; Han et al., 2024b; Wang et al., 2024b; Li et al., 2023; Zheng et al., 2024; Chiang et al., 2023; Du et al., 2024b). In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Prieto et al., 2016; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters et al., 2018; Wang et al., 2018; Yang et al., 2019; Miao et al., 2021), such as CNN (Kim, 2014), Bert (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on text classification tasks. However, the effectiveness of these NN-based approaches depends on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018). Furthermore, in real-world applications, such as Web Of Science (Kowsari et al., 2017) and IMDb (imd, [n.d.]), target labels for text often change based on the application’s requirements (Xia et al., 2021), leading to the dynamic few-shot text classification (DFSTC) task. As shown in Figure 1, the DFSTC task begins with a few initial classes (e.g., Machine Learning). As new topics like Computer Vision emerge in later rounds, models must adapt to these changes and accurately classify new examples with minimal labeled data. Hence, how to develop methods to dynamically classify text with limited labeled data available remains a necessity and open problem.
@@
Depending the technique for the DFSTC task, current models can be mainly categorized into two types, i.e., data augmentation-based models and RAG-based models. Firstly, data augmentation-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) create additional training data by mixing the pairs of the few-shot labeled data and assigning mixed labels indicating the validity for these created data based on the labels of each data pairs (Xia et al., 2021). Also, several data augmentation approaches create extra semantic-related content based on the label names (Meng et al., 2018; Meng et al., 2020) to enrich the training corpus further. However, due to the limited labeled data, and the label names may not always be available, the generated text data can have very limited patterns. Consequently, the text classification models trained on these generated data may over-fit on limited text data and are not generalizable (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023).
Figure 2. An overview of GraphRAG (Edge et al., 2024). It first extracts and indexes entities and relations as graph nodes and edges. Then it performs Top-K similarity-based retrieval for query-related subgraphs. Finally, it classifies texts based on the subgraph summary and the query text.
Recently, large language models (LLMs) (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), pre-trained on extensive corpora, have achieved significant success due to their superior and comprehensive text understanding abilities. However, researchers (Gao et al., 2023; Peng et al., 2024) discover that LLMs can struggle to understand their inputs due to a lack of task-specific knowledge. Therefore, the Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024) were proposed to combine the predicted text, candidate labels, and the retrieved side information (e.g., descriptions of text and labels, or documents) as LLM inputs, enable LLMs to understand the predicted text and labels more comprehensively with the help of side information from external sources. However, the incorporation of side information can further increase the input size and noise (Chen et al., 2023), which impedes the efficiency and effectiveness of LLMs (Sarthi et al., 2024).
@@
As a result, Compression-based RAG approaches (Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024) are proposed to compress the length of context and extract the key information. Specifically, LongLLMLingua (Xia et al., 2021) and CompaAct (Yoon et al., 2024) propose filtering uninformative texts based on the query. However, these approaches may result in information loss and degrade overall performance. Then, to address this, graph RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024; Peng et al., 2024; Zhang et al., 2025; Han et al., 2024a), such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2024), propose to to index the unstructured texts into the structured graphs. For example, in Figure 2, the basic idea of graph RAG considers text entities as graph nodes, and text sentences describing relations between these entity nodes as graph edges, later, it retrieves subgraphs base on the query text, and generates short summaries of these subgraphs as the side information for LLMs, reducing the input length for LLMs. However, existing Compression RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) still have three issues in the DFSTC task.
@@
    Uniform-indexing issue: Some of these approaches construct graphs by indexing extracted text chunks uniformly. They do not consider the varying importance and extraction confidence of each text chunk, which may provide incorrect and unreliable context for LLMs.
@@
    Threshold-dependent issue: These approaches select relevant information for each input text based on a globally predefined threshold. However, the optimal retrieval threshold can vary across different text samples; hence, relying on a globally predefined threshold can be suboptimal for the entire dataset.
@@
    Narrow-source issue: These approaches only retrieve side information from the each few-shot training text and ignore the important information among different the query texts, consequently limiting their retrieve accuracy and comprehensiveness.
To address these issues, we propose a novel hyperparameter-free Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot text classification, called GORAG. In general, GORAG constructs and maintains an adaptive online weighted graph by extracting side information from all target texts, and tailoring the graph retrieval for each input.
Firstly, GORAG constructs a weighted graph with keywords extracted from labeled texts as keyword nodes, and text labels as label nodes. The edge of this graph represents the relationship between keywords and labels. Secondly, GORAG employs an edge-weighting mechanism to assign different weights to edges created by the last step. An example of the constructed weighted graph is shown in Figure 3. The edge weights represent the keywords’ importance and relevance to the respective text’s label. Thirdly, GORAG constructs a minimum-cost spanning tree on the constructed weighted graph based on keywords of the query text, and then retrieves the label nodes within the spanning tree as candidate labels. Since the generated spanning tree is determined solely by the constructed graph and the keywords of the query text, GORAG achieves adaptive retrieval without relying on any human-defined retrieval thresholds, making the retrieval more precise regarding each query text. Lastly, GORAG applies an online indexing mechanism, which can enrich the weighted graph with the keywords from query texts, enhancing its few-shot learning performance by providing a more comprehensive retrieve source.
@@
  • We present a hyperparameter-free online RAG framework for DFSTC tasks, namely GORAG. The GORAG framework consists of three steps: graph construction, candidate label retrieval, and text classification, aim at addressing the uniform-index, threshold-dependent and narrow-source issues.
@@
  • To avoid any human-selected thresholds, we formulate the candidate label retrieval problem which is akin to the NP-hard Steiner Tree problem. To solve this problem, provide an efficient and effective solution to generate candidate labels.
@@ 2. Preliminary and Related Works
We first introduce the preliminaries of dynamic few-shot text classification in Section 2.1 and then discuss the related works for DFSTC task in Section 2.2. The important notations used in this paper are listed in Table 6 in Appendix A.
2.1. Dynamic Few-shot Text Classification Task
Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world application that involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$. Traditional approaches require fine-tuning with large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available. Recently, few-shot learning addresses this limitation by enabling models to classify text with only a small number of labeled examples per class (Meng et al., 2018; Meng et al., 2020). On the other hand, dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024d) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones. Combining these aspects, Dynamic Few-Shot Text Classification (DFSTC) task allows the model to handle evolving classification tasks with minimal labeled data, and better suited for real world scnerio (Xia et al., 2021).
In DFSTC task, the model is provided with multiple rounds of new class updates. Specifically, In each round $r$, a new set of classes $\mathcal{Y}_{new}^{r}$ is introduced, and the labelled dataset for $\mathcal{Y}_{new}^{r}$ is denoted as $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{new}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y_{i}\in\mathcal{Y}_{new}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$. Also, we denote candidate cumulative labels from the first round to the $r$-th round as $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}_{new}^{i}$. Formally, at the $r$-th round, given the candidate labels $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}_{new}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}={f}^{r}_{\theta}(t_{u},\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{u}$. Then, we can get the predicted label $y^{r*}\in\mathcal{Y}^{r}$ for the unseen text $t_{u}$ as follows.
(1) $y^{r*}=\arg\max_{y^{r}\in\mathcal{Y}^{r}}({f}_{\theta}^{r}(t_{u},\mathcal{Y}^{r})[y^{r}])$
@@ 2.1. Dynamic Few-shot Text Classification Task
2.2. Dynamic Few-shot Text Classification Model
Current DFSTC models can be broadly categorized into NN-based models and RAG-based models, RAG-based models can be further categorized into Long Context RAG models, and Compression-based RAG models.
@@ 2.2.1. NN-based models
NN-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) generate additional data contrastively based on the few-shot labeled data and the text formed label names to train the NN-based classifier. However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023). Also, as the text-formed label names are not always available in real-world scenarios, these NN-based models may not always be applicable.
Figure 3. An example of the constructed weighted graph, where red nodes denote the label nodes and black nodes denote the keyword nodes. The lower the edge cost is, the more related the keyword is to the respective label.
Figure 4. An overview of GORAG in each round. In Part 1, GORAG constructs a weighted graph based on keywords extracted from the few-shot training data. In Part 2, GORAG performs adaptive graph retrieval and outputs the candidate labels, which is a subset of the original target label set. In Part 3, GORAG first classifies texts into one of the candidate labels, and then applies online indexing to update the graph with newly extracted nodes from query text.

### v4

# GORAG: Graph-based Online Retrieval Augmented Generation for Dynamic Few-shot Social Media Text Classification

@@ Abstract.
Text classification is vital for Web for Good applications like hate speech and misinformation detection. However, traditional models (e.g., BERT) often fail in dynamic few-shot settings where labeled data are scarce, and target labels frequently evolve. While Large Language Models (LLMs) show promise in few-shot settings, their performance is often hindered by increased input size in dynamic evolving scenarios. To address these issues, we propose GORAG, a Graph-based Online Retrieval-Augmented Generation framework for dynamic few-shot text classification. GORAG constructs and maintains a weighted graph of keywords and text labels, representing their correlations as edges. To model these correlations, GORAG employs an edge weighting mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a tailored minimum-cost spanning tree for each input. Empirical evaluations show GORAG outperforms existing approaches by providing more comprehensive and precise contextual information. Our code is released at: https://github.com/Wyb0627/GORAG.
Large Language Model, Retrieval-Augmented Generation, Few-shot Text Classification, Hate Speech Detection, Social Media
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3793011isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Social networksccs: Information systems Clustering and classificationccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Figure 1. An example of DFSTC task with 2 rounds. In Round 1, the text is classified as Threatening Language. In Round 2, another text is classified as a new label: Identity Insults.
@@
Text content classification is a fundamental task for various social media applications, such as hate speech detection (Geissler et al., 2025; Hee et al., 2024; Chen et al., 2024b; Djuric et al., 2015; Subramanyam et al., 2024; Mandal et al., 2024; Cao et al., 2024; Ji et al., 2024; Guo et al., 2024), misinformation/disinformation detection (Luo et al., 2024; Nesterov et al., 2024; Liu et al., 2025; Toughrai et al., 2025; Cinus et al., 2025; Ashkinaze et al., 2024; Chen et al., 2022; Shang et al., 2024; Ye et al., 2024) , and suicide risk detection (Kang et al., 2024; Li et al., 2025; Naseem et al., 2025; Ge et al., 2025). In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters and others, 2018; Wang et al., 2018; Yang et al., 2019), such BERT (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on social media text content classification tasks. However, the effectiveness of these NN-based approaches (Johnson and Zhang, 2015; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters and others, 2018; Wang et al., 2018; Yang et al., 2019) depends on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018; Nguyen and Rudra, 2024). Furthermore, in real-world applications, such as Abuse Detection (Vidgen et al., 2021), new kinds of social media abuse can frequently appear with the changing needs of user analytics and progress in social research development, introducing the need to classify texts into new, extra labels (Xia et al., 2021), leading to the Dynamic Few-shot Social media Text Classification (DFSTC) task. As shown in Figure 1, the DFSTC task starts with an initial set of classes (e.g., Threatening Language). As new classes like Identity Insults emerge in later rounds, it becomes challenging to obtain extensive labeled data for these new categories. Models must therefore adapt to such changes and accurately classify new examples with limited labeled data. Developing methods to dynamically classify text with limited labeled data remains a crucial and open problem.
@@
Depending on the technique for the DFSTC task, current models can be mainly categorized into two types, i.e., data augmentation-based models and RAG-based models. Firstly, data augmentation-based models (Meng et al., 2018, 2020; Xia et al., 2021) create additional data for fine-tuning small language models (e.g., RoBERTa (Liu, 2019)), by mixing the pairs of the few-shot labeled data and assigning mixed labels indicating the validity for these created data based on the labels of each data pairs (Xia et al., 2021). However, due to the limited labeled data, and the label names may not always be available; furthermore, the additional data created can have very limited patterns. Consequently, the classification models trained with these created data may over-fit and are not generalizable (Liu et al., 2023, 2024a; Lei et al., 2023).
Recently, Large Language Models (LLMs) (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024a) have excelled in social media text classification through their superior understanding abilities. However, they often struggle with task-specific knowledge (Gao et al., 2023; Zhang et al., 2023) and lack accurate reasoning clues to explain responses (Peng et al., 2024; Wang et al., 2024b; Dong et al., 2025; Qian et al., 2025). In domains like hate speech or misinformation detection, these flaws can lead to undetected harmful content or wrongful penalties. Consequently, Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Xia et al., 2021; Yoon et al., 2024) were proposed to integrate retrieved side information (e.g., label descriptions) as external clues for more accurate classification. However, the incorporation of side information can further increase the input size (Chen et al., 2023), which often overwhelms critical details with noises. As the key clues of hate speech or misinformation are usually short sentences within long contexts, the ignorance of fine-grained, entity-level information can degrade LLM classification performance. Although proprietary LLMs (Hurst et al., 2024) can handle such extended inputs, they inevitably suffer from high token costs.
@@
To achieve finer-grained RAG, and better preserve the entity-level information, Graph-based RAG approaches (Edge et al., 2024; Guo et al., 2025; Jimenez Gutierrez et al., 2024; Peng et al., 2024; Zhang et al., 2025; Han et al., 2024; gutiérrez2025ragmemory), such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2025), propose to index the unstructured texts into the structured graphs. These approaches extract entities and relations from texts as graph nodes and edges, and their retrievals are based on the graph components mapped from the query text. However, existing Graph-based RAG approaches (Edge et al., 2024; Guo et al., 2025; Jimenez Gutierrez et al., 2024) still have three issues in the DFSTC task.
@@
    Uniform-indexing issue: Some of these approaches construct graphs by indexing extracted text chunks uniformly. They do not consider the varying importance of each text and extraction confidence of keyword entity. For example, in Figure 1, keyword lazy is related to all round 2 labels with different importance. Hence, uniform-indexing may provide inaccurate graph context for LLM classification.
@@
    Non-adaptive Retrieval issue: These approaches select relevant information for each input text based on a globally predefined threshold. However, the optimal retrieval threshold can vary across different query text and different rounds; hence, this non-adaptive retrieval approach can be suboptimal.
@@
    Narrow-source issue: These approaches only retrieve side information from a limited number of few-shot training text and ignore the important information among query texts. Since the new keywords related to certain labels in hate speech or misinformation detection emerge frequently in social media, this issue consequently limits models’ retrieval comprehensiveness.
To address these issues, we propose a novel Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot social media text classification, called GORAG. In general, GORAG constructs and maintains an adaptive online weighted graph by extracting side information from all target texts, and tailoring the graph retrieval for each input.
Firstly, GORAG constructs a weighted graph with keywords extracted from labeled texts as keyword nodes, and text labels as label nodes, to help the model better capture graph topology (Yang et al., 2024b; Zhao et al., 2024; Lin et al., 2024). To model relationships, GORAG employs an edge-weighting mechanism to assign different weights to edges, representing keywords’ importance and relevance to the respective text’s label. Secondly, GORAG constructs a minimum-cost spanning tree based on query keywords and retrieves the label nodes within it as candidate labels. Since this MST is determined solely by the constructed graph and the keywords of the query text, GORAG achieves adaptive retrieval without human-defined thresholds, ensuring precision for each input. Finally, GORAG applies an online indexing mechanism to enrich the graph with keywords from query texts. By calculating weights that reflect the semantic relevance of keywords to labels, GORAG effectively reduces the potential noise introduced by inaccurate LLM classification during real-time updates, ensuring the graph remains a robust and comprehensive retrieval source
@@
  • We present an online RAG framework for DFSTC tasks, namely GORAG, which consists of three steps: Graph Construction, Candidate Label Retrieval, and Text Classification, aiming to address the Uniform-indexing, Non-adaptive Retrieval, and Narrow-source issues.
@@
  • To achieve adaptive retrieval, we formulate the candidate label retrieval problem which is akin to the NP-hard Steiner Tree problem, and provide an efficient and effective solution based on the greedy algorithm.
  • Extensive experiments from the social media contextual abuse benchmark and the web news misinformation detection benchmark demonstrate the effectiveness and efficiency of GORAG.
@@ 2. Preliminary and Related Works
We first introduce the preliminaries of Dynamic Few-shot Social media Text Classification (DFSTC) in Section 2.1 and then discuss the related works for DFSTC task in Section 2.2. The important notations used in this paper are listed in Table 7 in Appendix A.
2.1. DFSTC Task
Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world web for good applications, including social media hate speech detection (Geissler et al., 2025; Hee et al., 2024; Chen et al., 2024b; Djuric et al., 2015; Subramanyam et al., 2024; Mandal et al., 2024; Cao et al., 2024; Ji et al., 2024; Guo et al., 2024), misinformation/disinformation detection (Luo et al., 2024; Nesterov et al., 2024; Liu et al., 2025; Toughrai et al., 2025; Cinus et al., 2025; Ashkinaze et al., 2024; Chen et al., 2022; Shang et al., 2024; Ye et al., 2024). It involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$. Traditional approaches require fine-tuning with large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available. Recently, few-shot learning addresses this limitation by enabling models to classify social media texts into various classes with only a small number of labeled examples per class (Meng et al., 2018, 2020). On the other hand, dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024a) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones, as frequently involve extensive human labeling for these new classes can be difficult. For example, in Figure 1, new and more fine-grained classes of hate speech, such as Identity Insults, can serve as complements to categories that already exists.
Combining these aspects, Dynamic Few-Shot Social media Text Classification (DFSTC) task allows the model to handle evolving classification tasks with minimal labeled data, and is better suited for real world scenarios (Xia et al., 2021). In DFSTC task, the model is provided with multiple rounds of new class updates. Specifically, in each round $r$, a set of new labels $\mathcal{Y}^{r}_{new}$ is introduced. The complete candidate label set for round $r$, denoted as $\mathcal{Y}^{r}$, is the union of all new labels from rounds $i=1$ to $r$: $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}^{i}_{new}$. In round $r$, the labeled training set for class labels in $\mathcal{Y}^{r}$ is denoted as $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y_{i}\in\mathcal{Y}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$. Formally, at the $r$-th round, given the labels in $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}={f}^{r}_{\theta}(t_{u},\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{u}$. Then, we can get the predicted label $y^{r*}\in\mathcal{Y}^{r}$ for the unseen text $t_{u}$ as follows.
(1) $y^{r*}=\arg\max_{y^{r}\in\cup_{i=1}^{r}\mathcal{Y}^{i}}({f}_{\theta}^{r}(t_{u},\cup_{i=1}^{r}\mathcal{Y}^{i})[y^{r}])$
@@ 2.1. DFSTC Task
2.2. DFSTC Models
Current Dynamic Few-Shot Social media Text Classification (DFSTC) models can be broadly categorized into NN-based models and RAG-based models, RAG-based models can be further categorized into Long Context RAG, and Graph-based RAG models.
@@ 2.2.1. NN-based models
NN-based models (Meng et al., 2018, 2020; Xia et al., 2021) generate additional data contrastively based on the few-shot labeled data and the text formed label names to train the NN-based classifier. However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023, 2024a; Lei et al., 2023). In web for good applications such as hate speech detection, different kinds of hate speech contents are usually a minority compared with neutral contents, this highly unbalanced data distribution can further overfit the NN-based models and affect their performance (Liu et al., 2023).
Figure 2. An overview of GORAG in each round. In Part 1, GORAG constructs a weighted graph based on keywords extracted from the few-shot training data. In Part 2, GORAG performs adaptive graph retrieval and outputs the candidate labels, which is a subset of the original target label set. In Part 3, GORAG first classifies texts into one of the candidate labels, and then applies online graph indexing to update the graph with query text extracted entities that not yet existed in the graph.